# General Order Virtual Element Methods for Neumann Boundary Optimal Control Problems in Saddle Point Formulation

### v1

@@ Acknowledgements
The authors acknowledge the fundings from INdAM - GNCS Projects (CUP: E53C23001670001). AB and FM acknowledge funding by the Italian Ministry of University and Research (MUR) through the project MUR-M4C2-1.1-PRIN 2022 CUP: E53D23005820006. AB ackowledges funding by the European Union through Next Generation EU, M4C2, PRIN 2022 PNRR project CUP: E53D23017950001 and PNRR M4C2 project of CN00000013 National Centre for HPC, Big Data and Quantum Computing (HPC) CUP: E13C22000990001. MS thanks the “20227K44ME - Full and Reduced order modelling of coupled systems: focus on non-matching methods and automatic learning (FaReX)” project – funded by European Union – Next Generation EU within the PRIN 2022 program (D.D. 104 - 02/02/2022 Ministero dell’Università e della Ricerca). This manuscript reflects only the authors’ views and opinions and the Ministry cannot be considered responsible for them.

### v2

@@ Acknowledgements
The four authors are members of the Gruppo Nazionale Calcolo Scientifico (GNCS) at Istituto Nazionale di Alta Matematica (INdAM). The authors kindly acknowledge financial support by INdAM-GNCS Project 2025 (CUP: E53C24001950001). AB and FM acknowledge funding by the Italian Ministry of University and Research (MUR) through the project MUR-M4C2-1.1-PRIN 2022 CUP: E53D23005820006. AB ackowledges funding by the European Union through Next Generation EU, M4C2, PRIN 2022 PNRR project CUP: E53D23017950001 and PNRR M4C2 project of CN00000013 National Centre for HPC, Big Data and Quantum Computing (HPC) CUP: E13C22000990001. MS thanks the “20227K44ME - Full and Reduced order modelling of coupled systems: focus on non-matching methods and automatic learning (FaReX)” project – funded by European Union – Next Generation EU within the PRIN 2022 program (D.D. 104 - 02/02/2022 Ministero dell’Università e della Ricerca). This manuscript reflects only the authors’ views and opinions and the Ministry cannot be considered responsible for them.